# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v3

@@ V Non-binary $2$-$b$-Burst-Deletion Correcting Codes
In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$. It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
Then $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. Like in Section IV-A, let $U(\mathbf{x})_{1,I_{j}}$ ($1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$) be all runs in $U(\mathbf{x})_{1}$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$.
Recall that $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$ We associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$. See Figure 1 for an illustration. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
@@
Lemma V.2
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.3. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
@@ 2. $(2)$
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.3, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.4, we can
@@
By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Note that for the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
In the following, we always assume $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Furthermore, suppose that $A(\mathbf{x})_{1}$ is $d$-regular. We need to find conditions on $\mathbf{x}$ which enable us to recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$. As in Section IV, we split our discussion into Sections V-A and V-B. Our main result in this section is presented in Theorem V.1.
V-A When Case (1) in Lemma V.2 happens
@@
We can assume $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma V.2.
Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$. Recall that $\left|I_{j}^{L}\right|\leq b-1$. So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N-1]$, where $N=q^{b-1}$. Note that $I_{1}^{L}=\emptyset$. For uniform discussion, we also let $\mathbf{x}_{I_{1}^{L}}=0$ and $\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}=\mathbf{x}_{\left[n-b+2,n\right]}$. Now define $L(\mathbf{x})=\mathbf{x}_{I_{1}^{L}}\cdots\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}}\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}\in[0,N-1]^{r\left(U(\mathbf{x})_{1}\right)+2}$. In other word, sequence $L(\mathbf{x})$ is the concatenation of $\mathbf{x}_{I_{1}^{L}},\ldots,\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}},\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}$. By Equation 14, the value of $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\notin\{j_{1},j_{1}+1,j_{2},j_{2}+1\}$.
Since $j_{2}\geq j_{1}+2$, it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures. To determine the values of $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, we apply Lemma II.7. Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7. Define
@@
$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
Since $r(U(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+2)+10\log N+4}=2^{\log n+O_{q,b}(1)}$. Now the following lemma follows from above analysis and Lemma II.7.

### v4

@@ 2. $(2)$
Notations: In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$. Denote the $i$-th row of $A(\mathbf{x})$ and $U(\mathbf{x})$ by $A(\mathbf{x})_{i}$ and $U(\mathbf{x})_{i}$, respectively. It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\},$
which is a subset of $[n]$. Then we have $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$. For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$. By definition, $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$, we associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$. See Figure 1 for an illustration. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
@@
Lemma IV.5
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.2. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can
@@
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.3, we can
@@
By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later. Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$. Then by II.2, we have $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Since $U(\mathbf{x})_{1}$ is regular, Lemma IV.2 ensures that when given $\eta\left(U(\mathbf{x})_{1}\right)$, one can uniquely recover $U(\mathbf{x})_{1}$ from $U\left(\mathbf{x}^{\prime}\right)_{1}$. Then by Lemma IV.3, we can know the approximate positions of the two deletions in $U(\mathbf{x})_{1}$. According to Lemma IV.5, we can approximately determine the positions of the two bursts in $\mathbf{x}$. Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C. Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7.
Before proceeding, we need the following observation.
Observation IV.1
Let $m>b$. Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}\in\mathcal{B}_{1}^{b}(\mathbf{u})$. Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
IV-B When Case (1) in Lemma IV.5 happens
@@
In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is know to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
@@
In the rest of this subsection, it is always assumed that $j_{2}-j_{1}\geq 2$, since otherwise, we have $\left|I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\cup I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\right|\leq 2bd\log(n/b)+b-1$, and therefore, this situation is covered by case (2) in Lemma IV.5 (or equivalently, case (2) in Lemma IV.3).
Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$. As a result, (13) implies $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$. Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}\cup I_{j_{1}}}$, $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
$\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}-1}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right),$ (14)
$\mathbf{x}^{\prime}_{\left[p_{j_{1}-1}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right).$
Similarly, we have
$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}-1}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right),$ (15)
$\mathbf{x}^{\prime}_{\left[(p_{j_{2}-1}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right).$
Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$. To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 00) and $\mathbf{x}^{\prime}_{\left[p_{-2}b+1,-b\right]}$ be an empty sequence.
By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion. Let $\phi(\cdot)$ be the function defined in Lemma II.2. Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $A(\mathbf{x})_{1}$ is $d$-regular, we have $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$. Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $[0,N-1]$, where $N=2^{\log\log(n)+O_{q,d,b}(1)}$. Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
Define $\overline{\phi}(\mathbf{x})=\phi\left(\mathbf{x}_{I_{0}^{\prime}}\right)\phi\left(\mathbf{x}_{I_{1}^{\prime}}\right)\cdots\phi\left(\mathbf{x}_{I_{r(U(\mathbf{x})_{1})}^{\prime}}\right)\in[0,N-1]^{r(U(\mathbf{x})_{1})+1}$. According to (13), when given $\mathbf{x}^{\prime}$, the value of $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures. To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.7 and define
$f(\mathbf{x})=\varphi\left(\overline{\phi}(\mathbf{x})\right).$
Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.7, $f(\mathbf{x})$ can be viewed as an integer in $[0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
The following lemma follows from the above analysis, Lemmas II.2 and II.7.
Lemma IV.6
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
IV-C When Case (2) in Lemma IV.5 happens
Recall that case (2) in Lemma IV.5 corresponds to case (2) in Lemma IV.3. Therefore, for case (2), instead of looking at $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we look at $A(\mathbf{x})$ and $A(\mathbf{x}^{\prime})$. In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$. Furthermore, by case (2) in Lemma IV.3, we found an interval $[c_{1},c_{2}]\subseteq[n/b]$ of length at most $3db\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$). By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c_{1}-1,c_{2}]$, for all $2\leq i\leq b$. Therefore, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$.
Note that $\left|[c_{1}-1,c_{2}]\right|$=$\left|J\right|+1\leq 3db\log(n/b)+1$. Let $P=3db\log(n/b)+1$ and define
Since $c_{1}$ and $c_{2}$ are known, we can find some $j_{0}$ such that $[c_{1}-1,c_{2}]\subseteq J_{j_{0}}$. Suppose $J_{j_{0}}=[d_{1},d_{2}]$. According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$. To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$.It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i,J_{j_{0}}}\right)$. Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$
Note that in the above definition of $g_{j}(\mathbf{x})$, we distinguish the two cases where $q>2$ and $q=2$. This is because when $q=2$, we have $A(\mathbf{x})=U(\mathbf{x})$, and hence the first row of $A(\mathbf{x})$ was already recovered.
Since $\left|J_{j}\right|\leq 2P=6db\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $+O_{q,d,b}(1)$. Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$). Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
$N_{2}=\begin{cases}2^{7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)},\mbox{ if }q>2,\\ 2^{7(b-1)\log\log n+O_{d,b}(1)},\mbox{ if }q=2.\\ \end{cases}$ (16)
For $s\in\{0,1\}$, let $L_{s}=\left\{1\leq j\leq\left\lceil n/bP\right\rceil-1:j\equiv s\pmod{2}\right\}$. For $s\in\{0,1\}$, define
$h^{(s)}(\mathbf{x})=\mathop{\sum}\limits_{j\in L_{s}}g_{j}(\mathbf{x})\pmod{N_{2}}.$ (17)
Lemma IV.7
For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for each $2\leq i\leq b$. In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
Proof:
Without loss of generality, assume $j_{0}\in L_{0}$. Then $g_{j}(\mathbf{x})$ is known to us for each $j\in L_{0}\setminus\{j_{0}\}$. By above discussion and Lemma II.4, to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ for all $2\leq i\leq b$, it is sufficient to obtain the value of $g_{j_{0}}(\mathbf{x})$. It follows from Equation 17, that
$g_{j_{0}}(\mathbf{x})\equiv\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}.$
Since $0\leq g_{j_{0}}(\mathbf{x})<N_{2}$, we have
$g_{j_{0}}(\mathbf{x})=\left(h^{(0)}(\mathbf{x})-\mathop{\sum}\limits_{j\in L_{0}\setminus\{j_{0}\}}g_{j}(\mathbf{x})\right)\pmod{N_{2}}.$
Now the proof is completed. ∎
We are ready to present our main result in this section.
Theorem IV.1
Suppose $q\geq 2$, $n>2b$ and $b>1$. Let $N_{1}$, $N_{2}$, $f(\mathbf{x})$, $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$ be defined as above. Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. For any $0\leq a<N_{0}$, $0\leq b<N_{1}$ and $0\leq c_{0},c_{1}<N_{2}$, define the code $\mathcal{C}_{1}$ as
@@
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }d\text{-regular},\\ \eta(U(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ h^{(0)}(\mathbf{x})=c_{0},h^{(1)}(\mathbf{x})=c_{1}\end{array}\right\}.$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
  • $q>2$ is even, $d=7$ and $n\geq 9$;
If $q=2$, $d=7$ and $n\geq 9$, the redundancy is upper bounded by $5\log n+14b\log\log n+O_{b}(1)$.
Proof:
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma IV.7 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{1}$ can correct two $b$-burst-deletions.
By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular. It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, and $c_{1}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{q^{n-1}}{N_{0}\cdot N_{1}\cdot(N_{2})^{2}},$
which implies by definition that
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+2\log N_{2}+1$
$=5\log n+14\log\log n+2\log N_{2}+O_{q,b}(1).$
Now the claim for redundancy follows from (16). ∎
V Non-binary Two-deletion Correcting Codes
In this section, it is always assumed that $q>2$. We aim to give a new construction of $q$-ary two-deletion correcting codes. Recall the definition of $u(\mathbf{x})$ for any $\mathbf{x}\in\Sigma_{q}^{n}$. If $\mathbf{x}^{\prime}\in\mathcal{B}_{2}(\mathbf{x})$, then $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}(u(\mathbf{x}))$. So if $u(\mathbf{x})$ belongs to a binary two-deletion correcting code, we can recover $u(\mathbf{x})$ from $u(\mathbf{x}^{\prime})$. We further assume that $\mathbf{x}$ (and thus $u(\mathbf{x})$) is $d$-regular. Then Lemma IV.3 asserts that we can either find two runs of $u(\mathbf{x})$ and each of them suffers one deletion, or find a substring of $u(\mathbf{x})$ of short length which suffers two deletions. This enables us to approximately determine error positions in $\mathbf{x}$. For the latter case, we need a $q$-ary two-deletion correcting code, which is given in Lemma II.4. For the former case, we need a $q$-ary single-deletion correcting code. We use the code given in [23]